If my badminton partner hits kill shots with a normal distributed mean of 90.4, and a standard deviation of 5.3, find the probability that he will really bomb in the next match, hitting with a kill shot lower than a 76.
Answer by stanbon(75887)

If my badminton partner hits kill shots with a normal distributed mean of 90.4, and a standard deviation of 5.3, find the probability that he will really bomb in the next match, hitting with a kill shot lower than a 76.
------
z(76) = (76-90.4)/5.3 = -2.7170
---
P(x < 76) = P(z < -2.7170) = 0.0033
================
Cheers,
Stan H.
================